sum of two numbers is 36.
One number is 3 times the other number.
What are the numbers?

Answer by feliz1965(151)

small number = x
bigger number = 3x
Their sum = x + 3x
So, x + 3x = 36
4x = 36
x = 36/4
x = 9
The small number is 9 and the bigger number is 3 times 9 or 27.